OpenStudy (karatechopper):

2 number cubes rolled 150 times find probability if the sum less than 4, 15 times

OpenStudy (anonymous):

I am not sure I understand it completely. You want the probability of rolling 2 dice 150 times and getting exactly 15 times 1-1 or 1-2?

OpenStudy (karatechopper):

i didn't understand it either i started off by doing 150 times 12

OpenStudy (anonymous):

I have think about this but still I am not sure what probability you need

OpenStudy (karatechopper):

experimental probability

OpenStudy (amistre64):

im guessing here; but the probability of rolling a sum of less than 4 on one roll is 3/36;

OpenStudy (amistre64):

times that by 15 and divide by 150?

OpenStudy (anonymous):

experimental probability is something different as I know. It is something like this example: Sam rolled a number cube 50 times. A 3 appeared 10 times. Then the experimental probability of rolling a 3 is 10 out of 50 or 20%.

OpenStudy (karatechopper):

well i looked up experimental probability, and then i did 15/150, reduced it 1/10, 10%

OpenStudy (karatechopper):

tht right?

OpenStudy (anonymous):

I am not sure

OpenStudy (amistre64):

could we make it binomial? p=3/36 , q=33/36, n=150, x=15?

OpenStudy (amistre64):

.08344 if we can :)

OpenStudy (karatechopper):

i did not understand tht

OpenStudy (amistre64):

if we can do a binomial probability; it would make it easier, but i dont recall the rules that well

OpenStudy (amistre64):

fixed number of trials; same probability of success, and some other things

OpenStudy (amistre64):

150 C 12 * p^(15) * q^(....) something

OpenStudy (amistre64):

that 12 should be a 15 ....